[Hallc_running] [New Logentry] RC Daily Update

dhamilto at jlab.org dhamilto at jlab.org
Thu Nov 29 11:05:01 EST 2018

Logentry Text:

- There were several issues through Swing and Owl as CW beam was restored to the hall, including a stuck harp, the wrong raster initialization and a stuck HMS collimator motion controller.
- Production data-taking on LH2 elastic singles started around 2am, and data-taking for all five settings (with dummy) were completed by the shift change this morning.
- We have reconfigured for LH2 elastic coincidence measurements, although beam is currently not available for two hours due to a Wien angle change for Hall B. It should be back at 11am.
- Mark took the opportunity to go to the hall and look at the HMS collimator and check it is ok (and not stuck somewhere as a result of the collimator motion problem last night). His recommendation is to postpone optics data-taking until next week following a planned reboot of the collimator motion controller during the beam studies / system recovery on Monday Day shift.
- When the beam is restored this morning, the plan is to take the LH2 elastic coincidence runs and move the spectrometers to their production settings.
- The crew chief just confirmed they are ready to do the SHMS fringe field beam steering study and harp energy scan as soon as we are done with the elastics.
- If all goes well, the plan would be to start production data-taking around end of day / beginning of owl for 80 hours.

RC Meeting:
- There will be an RC meeting today at 3:30pm, second floor counting house.

Notices in effect for shift crews:
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''
- Shift leaders, please fill out the check list once per shift!
- Shift leaders, please cross-check the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.

Planned accelerator configuration changes:
- None

Planned accesses to the Hall:
- Monday (during beam studies / system recovery): Mark and Ethan plan to reboot the collimator motion controller.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3632934
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181129/b456aea4/attachment-0001.html>

More information about the Hallc_running mailing list